A jet of water squirts out horizontally from a hole near the bottom of the tank . The tank is on a platform that is 1m above the ground .If the hole has a diameter of 3.5mm what is the height of the water level in the tank .

1
Expert's answer
2021-07-20T15:08:35-0400

Gives y=1m

Daimeter=3.5mm

Assume x=0.546m


"t^2=\frac{2y}{g}=\frac{2\times1}{9.8}=0.205sec"

"t=\sqrt{0.205}=0.452sec"


"v=\frac{x}{t}=\frac{0.546}{0.452}=1.20m/sec"

We have

"\frac{1}{2}\rho v^2=\rho gh"

"h=\frac{v^2}{2g}=\frac{(1.20)^2}{2\times9.8}=0.073cm"